A person hits bullets at 2m/s costant velocity at the rate of n per second, the mass of each bullet is m. Calculate the force.
a.2nm/t
b.2nm
c.nm/2
d.nm

Grade:12th pass

2 Answers

Arun
4 years ago
Dear student
 
As we know, force = (mass/ sec) * velocity = mn * 2 = 2mn
 
 
Hence option B is correct.
Hope it helps
 
Regards
Arun
Vikas TU
4 years ago
Dear student 
Force = mass * acceleration 
acceleration = velocity / time 
In 1 sec the force exerted will be 
Force = m * v/1 * n 
=> 2mn 
B is correct option